What is the quadtratic equation when the roots are 1+rad3 and 1-rad3
Answer by FrankM(1040)

(X-(1+sqrt(3)))(X-(1-sqrt(3)))=0
((X-1)-sqrt(3))((X-1)+sqrt(3))=0
(X-1)^2-(sqrt(3))^2=0
X^2-2X+1-3=0
X^2-2X-2=0